# Path-based packing of icosahedral shells into multi-component aggregates

### v1

@@
Multi-component aggregates are being intensively researched in various fields because of their highly tunable properties and wide applications. Due to the complex configurational space of these systems, research would greatly benefit from a general theoretical framework for the prediction of stable structures, which, however, is largely incomplete at present. Here we propose a general theory for the construction of multi-component icosahedral structures by assembling concentric shells of different chiral and achiral types, consisting of particles of different sizes. By mapping shell sequences into paths in the hexagonal lattice, we establish simple and general rules for building a wide variety of magic icosahedral structures, and we evaluate the optimal size-mismatch between particles in the different shells. Our predictions are confirmed by numerical simulations for different systems.
@@ Introduction
The wide range of possible compositions and the extremely rich configurational space of multi-component systems are key to their broad success. However, such inherent complexity poses major challenges to the design and synthesis of nanoaggregates with well-defined and durable configurations. A general theory for the prediction of stable multicomponent structures would be of great help, but there is currently a lack of results in this regard.
@@
Here we show that it is possible to establish general criteria for assembling multiple components into highly symmetrical geometric structures. We consider building blocks (hereafter referred to as particles) of different sizes, assembling them into multi-shell aggregates of icosahedral symmetry. The icosahedron is chosen because combines the maximum symmetry with the most compact shape. These properties are advantageous both for inorganic [13, 14] and for biological [15, 16, 17] systems. Accordingly, achiral and chiral icosahedral structures have been observed in clusters and nanoparticles [18, 19, 20, 21, 22], colloidal aggregates [23, 24, 25], intermetallic compounds and quasicrystals [26, 27, 28], viruses, bacterial organelles, DNA and protein aggregates [29, 30, 31, 32, 33]. Moreover, due to the contraction of pair distances between particles in adjacent shells [34], the icosahedron is naturally suitable for accommodating particles of different sizes in different shells.
In our approach, we generalize and unify concepts from crystallography [34] and structural biology [15, 17] to propose a general method for assembling multiple achiral and chiral shells into multi-component icosahedra. The key point is the mapping of multi-shell structures into paths in the hexagonal lattice, which naturally leads to a design principle of multi-component aggregates corresponding to new series of icosahedral magic numbers. The aggregates are stabilized by the size mismatch between particles of different shells. This design principle is validated by numerical calculations for several model systems, and by ab initio calculations for alkali metal clusters. In addition, growth simulations of alkali and transition metal nanoaggregates show that atoms naturally self-assemble into the same multi-shell icosahedral structures predicted by our theory.
@@
The starting point of our design strategy is the well-known theory of Caspar and Klug [15, 35, 17], originally proposed for rationalizing and predicting the architecture of icosahedral viral capsids. Specifically, Caspar and Klug (CK) proposed a general method for the construction of individual icosahedral shells, which is based on cutting and folding leaflets from the two-dimensional hexagonal lattice and produces achiral and chiral arrangements of particles on the icosahedral surface.
The CK construction is shown in Fig. 1A. A segment is drawn between points of coordinates (0,0) and $(h,k)$ with respect to the basis vectors of the hexagonal lattice; $h$ and $k$ are integer non-negative numbers, so that the segment always connects two lattice points. The segment is the base of an equilateral triangle, which is replicated 20 times to form a leaflet, which is then cut and folded to generate an icosahedral shell with a well-defined surface lattice.
In the CK theory, the triangulation number $T$ of an icosahedral shell is defined as the square length of the triangular edge, and is calculated as $T=h^{2}+k^{2}+hk$. Edge and radius of the shell are $\sqrt{T}$ and $\sin(2\pi/5)\sqrt{T}$, respectively. Assigning one particle to each lattice point, the shell contains $210\,T+$ 2 particles (see section S1 of the supplementary materials).
@@
The main point of our construction consists in assembling concentric shells into aggregates by drawing paths in the hexagonal lattice. As a first example, we consider a path along a coordinate axis, e.g., the $k$ axis (Fig. 2A), starting from $k=0$ and making steps $(0,k)\to(0,k+1)$ up to $k=i-1$. This path assembles $i$ concentric MC shells of larger and larger size into a Mackay icosahedron [34, 13], a well-known structure observed in many experiments on clusters [18, 37, 38, 19, 22], which is thus recovered as a special case of our construction. Mackay icosahedra are tightly packed structures consisting of 20 distorted tetrahedra [13] in which the particles are arranged according to the face-centered-cubic (fcc) lattice. The numbers of particles in a MC icosahedron made of $i$ shells is $N_{i}=(10i^{3}-15i^{2}+11i-3)/3$, which gives the series of magic numbers 1, 13, 55, 147, 309,…
In general, paths can be drawn by allowing different increments at each step. Here we deal with the simplest generalization, which consists of choosing between $(h,k)\to(h,k+1)$ and $(h,k)\to(h+1,k)$. In this case, the number of shells in a path from the origin to a point $(h,k)$ is $i=h+k+1$. Such paths are inspired by the Mackay path of Fig. 2A, in which one index is incremented at steps of one, whereas the other is kept constant and equal to zero. Allowing for different elementary steps in the hexagonal plane allows to build a wide variety of icosahedral structures, which retain the densely-packed character of the Mackay icosahedron.
We distinguish three cases, as shown in Fig. 2B. If $(h,k)$ is above the diagonal, the elementary move $(h,k)\to(h,k+1)$ conserves the chirality class of the shell, i.e. the shell we are adding belongs to the same class of the previous one. On the contrary, by the move $(h,k)\to(h+1,k)$ the chirality class is incremented. If $(h,k)$ is below the diagonal, the opposite applies. For $(h,k)$ in the diagonal, i.e. for shells of BG type, both steps conserve the class; specifically, the two steps are equivalent since the corresponding added shells are enantiomers, with the same size and particle arrangement but opposite chirality.
@@
In the example of Fig. 2, C and D, the core is made of five MC shells, i.e. it is terminated by the (0,4) shell. According to our path rules, the outer shell can be either the (0,5) MC shell (same class) or the (1,4) Ch1 shell (different class). We consider a simple model, in which all particles interact by the well-known Lennard-Jones (LJ) potential; specifically, the interaction energy is the same for all particles, the only difference being the equilibrium distance of pair interactions, which accounts for particles of different sizes (see the Methods section in the supplementary materials). In Fig. 2C, we calculate the binding energy depending on the size-mismatch, and determine the mismatch that minimizes the energy of the whole aggregate. For both MC and Ch1 shells this optimal mismatch is positive, i.e. it is favourable to have bigger particles in the outer shell than in the core, and it is larger for Ch1 than for MC shells. The optimal mismatch can be estimated also by geometric packing arguments (see section S1.4 in the supplementary materials) as
$sm_{i,i+1}=\frac{2\sin(2\pi/5)(1+\xi)\sqrt{T_{i}}+1}{2\sin(2\pi/5)\sqrt{T_{i+1}}-1}-1,$ (1)
@@
where $T_{i}$ and $T_{i+1}$ are triangulation numbers of shell $i$ and $i+1$, and $\xi\ll 1$ is an expansion coefficient of pair distances in the core, which depends on the number and the type of shells of the same component in the core. In Fig. 2, E and F the results of Eq. 1 are compared to those for LJ and Morse clusters. Data in Fig. 2E are obtained as in Fig. 2C, but for different sizes of the Mackay core, whereas in Fig. 2F we consider a cluster made of a Mackay core plus a Ch1 shell, to which we add a further shell of either Ch1 or Ch2 type. In all cases the agreement is good. Eq. 1 is thus a reliable guide for the semi-quantitative evaluation of the optimal mismatch, which demonstrates the key role of geometric factors in determining it.
@@
For the path of Fig. 2, the optimal mismatch is always positive, i.e. icosahedral aggregates benefit from having bigger particles in the outer shells; in addition, much larger mismatches are found for class-changing than for class-conserving steps. These features are general for all icosahedra constructed according to the path rules identified so far, and are key to design stable icosahedral aggregates and to predict their natural growth modes.
@@
Here we develop a path-based design strategy for icosahedral aggregates. First, a path is drawn according to the rules in Fig. 2 and, for each step, the optimal mismatch is estimated by Eq. 1. Then, particles of the appropriate sizes are associated with each shell.
An example is shown in Fig. 3, where the path connecting BG shells through neighbouring chiral shells is considered (Fig. 3A). This path alternates class-conserving and class-changing steps, spontaneously breaking mirror symmetries after the first BG shell. It produces a new series of magic numbers
@@
with $b=4$ and $b=9$ for even and odd $i$, i.e. $N_{i}=1,13,45,117,239,431..$ (see section S1.2.1 of the supplementary material).
@@
The clusters in Fig. 3C correspond to $i=3$ (Na13@K32, Na@${}_{13}@$Rb32) and to $i=4$ (Na@${}_{13}@$K@${}_{32}@$K72, Na@${}_{13}@$Rb@${}_{32}@$Rb72, Na@${}_{13}@$K@${}_{32}@$Rb72). In these clusters, the atomic species is always changed in class-changing steps, where a large mismatch is required, while it is changed or not in class-conserving steps, where the optimal mismatch is moderate, and therefore zero mismatch is expected to be acceptable. The stability of these clusters is verified by Density Functional Theory (DFT) calculations (data are reported in tables S1-S3), specifically by checking that exchanges of pairs of different atoms in adjacent shells produce energy increases. In contrast, if MC shells of these species are assembled in the same order, the resulting clusters are energetically unstable with respect to exchanges of atomic pairs, because the mismatch between species is too large for MC shells.

### v2

@@
Multi-component aggregates are being intensively researched in various fields because of their highly tunable properties and wide applications. Due to the complex configurational space of these systems, research would greatly benefit from a general theoretical framework for the prediction of stable structures, which, however, is largely incomplete at present. Here we propose a general theory for the construction of multi-component icosahedral structures by assembling concentric shells of different chiral and achiral types, consisting of particles of different sizes. By mapping shell sequences into paths in the hexagonal lattice, we establish simple and general rules for designing a wide variety of magic icosahedral structures, and we evaluate the optimal size-mismatch between particles in the different shells. The predictions of our design strategy are confirmed by molecular dynamics simulations and density functional theory calculations for several multi-component atomic clusters and nanoparticles.
@@ Introduction
The wide range of possible compositions and the extremely rich configurational space of multi-component systems are key to their broad success. However, such inherent complexity poses major challenges to the design and synthesis of nanoaggregates with well-defined and durable configurations. A general theory for the prediction of stable multicomponent structures would be an essential reference for the design and synthesis of nanoparticles for customised applications. However, such a theory is lacking at present.
Here we propose a theoretical approach which generalizes and unify concepts from crystallography [13] and structural biology [14, 15, 16] to develop a design strategy of multi-component clusters and nanoparticles.
@@
We consider multi-component aggregates formed by particles of different sizes, and establish general criteria for assembling these particles into highly symmetric multi-shell structures. In particular, we apply our approach to icosahedral structures. Icosahedra combine the maximum symmetry with the most compact shape. These properties favour energy stability in clusters and nanoparticles [17, 18] and give an evolutionary advantage in biological systems such as viruses [14, 16, 19]. Accordingly, icosahedra have been observed in a huge variety of systems, including clusters and nanoparticles [20, 21, 22, 23, 24], colloidal aggregates [25, 26, 27], intermetallic compounds and quasicrystals [28, 29, 30], viral capsids, bacterial organelles, DNA and protein aggregates [31, 32, 33, 34, 35].
Several research efforts can be found in the literature, starting with the seminal work of Caspar and Klug [14], dedicated to rationalising the structure of individual icosahedral shells, especially in the field of virus biology [15, 16, 36, 37]. Icosahedral shells are made of one layer of particles, which can be arranged according to achiral or chiral symmetries. Chiral shells present all rotational symmetry operations of the icosahedron, but lack its reflection planes.
On the contrary, there is no general theory for assembling together multiple shells of different radii, with achiral and chiral symmetries, into concentric arrangements to generate compact icosahedral aggregates. In metal nanoparticles and clusters, and in aggregates of colloidal particles, compact structures are much more commonly observed than single shells [20, 17, 21, 24, 25, 26, 27]. The theoretical efforts to assemble multi-shell icosahedra are rather limited. They began with the historical works of Bergman et al. [38] and Mackay [13], both concerning achiral structures only. In particular, Mackay constructed icosahedra by packing spheres of equal size arranged in shells around a central particle. Mackay icosahedra turns out to be built of 20 distorted tetrahedra, in which particle layers are arranged into the stacking of the face-centered cubic lattice (ABCABC…). Most icosahedra observed in metal clusters and nanoparticles are of Mackay type [20, 21, 22, 24]. Mackay also proposed a possible termination by a single shell in hexagonal close-packed (hcp) stacking (giving for example the sequence ABCABCB), known as the anti-Mackay shell. Anti-Mackay terminations have been recently observed in colloidal aggregates [25, 26]. Even fewer theoretical efforts have been devoted to construct icosahedra that include chiral shells. These efforts have been limited to adding a single chiral shell on top of an achiral Mackay core [39, 40, 41].
The key point of our theoretical approach is the mapping of multi-shell structures into paths in the hexagonal lattice. The mapping naturally leads to the design of new multi-shell sequences corresponding to new series of icosahedral magic numbers. Furthermore, the mapping allows us to predict which sequences exhibit spontaneous symmetry breaking from achiral to chiral structures. In the field of metal clusters and nanoparticles, chiral icosahedra are especially interesting for their applications to catalysis [41]. In all cases, we demonstrate that icosahedra are stabilized by the size mismatch between particles of different shells and evaluate the optimal mismatch for energetic stability. We note that, due to the contraction of pair distances between particles in adjacent shells [13], the icosahedron is naturally suitable for accommodating particles of different sizes in different shells. Our approach is generalized also to anti-Mackay shells, so that it allows us to derive a design principle for the structures observed in colloidal aggregates [25, 26, 27], and to predict new ones.
Our design strategy is validated by numerical calculations for several model systems and by ab initio calculations for alkali metal clusters. The predictions of our theory are further confirmed by simulations of the growth of alkali and transition metal nanoparticles including up to four different elements. These simulations show that atoms naturally self-assemble into the predicted multi-shell icosahedral structures, including those with symmetry-breaking shell sequences. Although the applications presented below concern atomic clusters and nanoparticles, due to the general character of its basic assumptions, our theory can be applied to the design of aggregates of other particle types, such aggregates of colloidal particles and complex molecules of biological interest.
@@
Here we develop our theoretical framework, whose starting point of is the well-known approach of Caspar and Klug [14, 16, 15], which was originally proposed for rationalizing and predicting the architecture of icosahedral viral capsids. Specifically, Caspar and Klug (CK) developed a general method for the construction of individual icosahedral shells, which is based on cutting and folding leaflets from the two-dimensional hexagonal lattice and produces achiral and chiral arrangements of particles on the icosahedral surface.
[…]
The CK construction is shown in Figure 1a. A segment is drawn between points of coordinates $(0,0)$ and $(h,k)$ with respect to the basis vectors of the hexagonal lattice; $h$ and $k$ are integer non-negative numbers, so that the segment always connects two lattice points. The segment is the base of an equilateral triangle, which is replicated 20 times to form a leaflet, which is then cut and folded to generate an icosahedral shell with a well-defined surface lattice.
In the CK theory, the triangulation number $T$ of an icosahedral shell is defined as the square length of the triangular edge, and is calculated as $T=h^{2}+k^{2}+hk$. Edge and radius of the shell are $\sqrt{T}$ and $\sin(2\pi/5)\sqrt{T}$, respectively. Assigning one particle to each lattice point, the shell contains $210\,T+$ 2 particles (see Supplementary Note 1.1).
@@
Figure 2: Mapping icosahedra into paths. a Path of a Mackay icosahedron: $i=5$ shells are assembled on top of each other. b The three possible cases when choosing between $(h,k)\to(h,k+1)$ and $(h,k)\to(h+1,k)$ at each step of the path.
The main point of our construction consists in assembling concentric shells into aggregates by drawing paths in the hexagonal lattice. As a first example, we consider a path along a coordinate axis, e.g., the $k$ axis (Figure 2a), starting from $k=0$ and making steps $(0,k)\to(0,k+1)$ up to $k=i-1$. This path assembles $i$ concentric MC shells of larger and larger size into a Mackay icosahedron [13, 17], a well-known structure observed in many experiments on clusters [20, 21, 24, 42, 43], which is thus recovered as a special case of our construction. Mackay icosahedra are tightly packed structures consisting of 20 distorted tetrahedra [17] in which the particles are arranged according to the face-centered-cubic (fcc) lattice. The numbers of particles in a MC icosahedron made of $i$ shells is $N_{i}=(10i^{3}-15i^{2}+11i-3)/3$, which gives the series of magic numbers $1$, $13$, $55$, $147$, $309$,…
In general, paths can be drawn by allowing different increments at each step. Here we deal with the simplest generalization, which consists of choosing between $(h,k)\to(h,k+1)$ and $(h,k)\to(h+1,k)$. In this case, the number of shells in a path from the origin to a point $(h,k)$ is $i=h+k+1$. Such paths are inspired by the Mackay path of Figure 2a, in which one index is incremented at steps of one, whereas the other is kept constant and equal to zero. Allowing for different elementary steps in the hexagonal plane allows to build a wide variety of icosahedral structures, which retain the densely-packed character of the Mackay icosahedron.
We distinguish three cases, as shown in Figure 2b. If $(h,k)$ is above the diagonal, the elementary move $(h,k)\to(h,k+1)$ conserves the chirality class of the shell, i.e. the shell we are adding belongs to the same class of the previous one. On the contrary, by the move $(h,k)\to(h+1,k)$ the chirality class is incremented. If $(h,k)$ is below the diagonal, the opposite applies. For $(h,k)$ in the diagonal, i.e. for shells of BG type, both steps conserve the class; specifically, the two steps are equivalent since the corresponding added shells are enantiomers, with the same size and particle arrangement but opposite chirality.
@@
Figure 3: Optimal size mismatch. a Binding energy per atom for clusters made of a core with five MC shells plus a sixth shell, which is either MC or Ch1. Interactions are of Lennard-Jones (LJ) type (see Methods Section). The particles of the sixth shell differ from those in the core only by their size. The energy is in units of the $\varepsilon$ of the LJ potential. The structures are shown in b. c Comparison of the optimal mismatch between steps $(0,k)\to(0,k+1)$ and $(0,k)\to(1,k)$, corresponding to additional MC and Ch1 shells on an MC core, respectively. The optimal mismatch of Eq. 3 (GEO values) is compared to Lennard-Jones (LJ) and Morse potential results. The Morse potential data are given for three values (4,5,6) of the parameter $\alpha$, that regulates the width of the potential well (see Methods Section). d Comparison between steps $(1,k)\to(1,k+1)$ and $(1,k)\to(2,k)$, corresponding to additional Ch1 and Ch2 shells on a core containing $k$ MC and one Ch1 shells, respectively. The optimal mismatch of Eq. 3 (GEO) is compared to LJ data
@@
In the example of Figure 3a-b, the core is made of five MC shells, i.e. it is terminated by the $(0,4)$ shell. According to our path rules, the outer shell can be either the $(0,5)$ MC shell (same class) or the $(1,4)$ Ch1 shell (different class). We consider a simple model, in which all particles interact by the well-known Lennard-Jones (LJ) potential; specifically, the interaction energy is the same for all particles, the only difference being the equilibrium distance of pair interactions, which accounts for particles of different sizes (see the Methods section in the supplementary materials). In Figure 2a, we calculate the binding energy depending on the size mismatch, and determine the mismatch that minimizes the energy of the whole aggregate. For both MC and Ch1 shells this optimal mismatch is positive, i.e. it is favourable to have bigger particles in the outer shell than in the core, and it is larger for Ch1 than for MC shells.
The optimal mismatch can be estimated also by geometric packing arguments, which are discussed in details in Supplementary Note 1.4. Here we recall only the main assumption and give the final result. In order to evaluate the optimal mismatch between particles of shells $i$ and $i+1$ it is reasonable to impose that
$r_{i+1}-r_{i}=\frac{d_{i}+d_{i+1}}{2}$ (1)
where $r_{i}$ and $r_{i+1}$ are the radii of the respective shells, $d_{i}$ and $d_{i+1}$ are the particle diameters in these shells. For atomic systems, these diameters are estimated from the nearest-neighbour distance of atoms in their crystal lattice. Recalling that the relation between the radii and the triangulation numbers of the shells is
$r_{i}=d_{i}\sin\left(\frac{2\pi}{5}\right)\sqrt{T_{i}}$ (2)
one finally obtains the following approximate expression for the optimal size mismatch
$sm_{i,i+1}=\frac{2\sin\left(\frac{2\pi}{5}\right)(1+\xi)\sqrt{T_{i}}+1}{2\sin\left(\frac{2\pi}{5}\right)\sqrt{T_{i+1}}-1}-1.$ (3)
Here $\xi\ll 1$ is an expansion coefficient of pair distances in the core, which depends on the number and the type of shells of the same component in the core (see Supplementary Note 1.4).
@@
In Figure 3c-d the results of Eq. 1 are compared to those for LJ and Morse clusters. Data in Figure 3c are obtained as in Figure 3a, but for different sizes of the Mackay core, whereas in Figure 3d we consider a cluster made of a Mackay core plus a Ch1 shell, to which we add a further shell of either Ch1 or Ch2 type. In all cases the agreement is good. Eq. 1 is thus a reliable guide for the semi-quantitative evaluation of the optimal mismatch, which demonstrates the key role of geometric factors in determining it.
@@
For the paths of Figure 3, the optimal mismatch is always positive, i.e. icosahedral aggregates benefit from having bigger particles in the outer shells; in addition, much larger mismatches are found for class-changing than for class-conserving steps. These features are general for all icosahedra constructed according to the path rules identified so far, and are key to design stable icosahedral aggregates and to predict their natural growth modes.
@@
Our theory is now applied to the path-based design of icosahedral aggregates. In order to construct a multi-shell icosahedral aggregate, a path is drawn according to the rules in Figure 2 and, for each step, the optimal mismatch is estimated by Eq. 1. Then, particles of the appropriate sizes are associated with each shell.
Figure 4: A path with spontaneous symmetry breaking. a A path connecting all BG shells through their nearby chiral shells. Black and grey arrows indicate paths with right and left-handed chiral shells. A generic path of this type may alternate both chiralities, corresponding to combinations of black and grey arrows. Shells are enumerated along the black-arrow path. b Optimal mismatch $sm_{i,i+1}$ according to Eq. 1 for the path in a. Points connected by brown and green lines correspond to changes of species every shell and every two shells. The optimal $sm_{i,i+1}$ are compared to the mismatch between pairs of atomic species, indicated by squares and diamonds. The triangles correspond to the optimal mismatch between two MC shells and two Ch1 shells. c Structures for $i=2,3,4$ along the path of a, with their compositions related to specific systems.
An example is shown in Figure 4, where the path connecting BG shells through neighbouring chiral shells is considered (Figure 4a). This path alternates class-conserving and class-changing steps, spontaneously breaking mirror symmetries after the first BG shell. It produces a new series of magic numbers
@@
with $b=4$ and $b=9$ for even and odd $i$, i.e. $N_{i}=1,13,45,117,239,431..$ (see Supplementary Note 1.2).
@@
The clusters in Figure 4c correspond to $i=3$ (Na13@K32, Na@${}_{13}@$Rb32) and to $i=4$ (Na@${}_{13}@$K@${}_{32}@$K72, Na@${}_{13}@$Rb@${}_{32}@$Rb72, Na@${}_{13}@$K@${}_{32}@$Rb72). In these clusters, the atomic species is always changed in class-changing steps, where a large mismatch is required, while it is changed or not in class-conserving steps, where the optimal mismatch is moderate, and therefore zero mismatch is expected to be acceptable. The stability of these clusters is verified by Density Functional Theory (DFT) calculations (data are reported in Table 1 and Supplementary Tables 1-3), specifically by checking that exchanges of pairs of different atoms in adjacent shells produce energy increases. In contrast, if MC shells of these species are assembled in the same order, the resulting clusters are energetically unstable with respect to exchanges of atomic pairs, because the mismatch between species is too large for MC shells. Our theory therefore provides new possibilities for constructing icosahedra in systems where the traditional Mackay icosahedra would not be stable.
3 SHELLS 4 SHELLS
MC@BG MC@MC MC@BG@Ch1 MC@MC@MC
Na@K +0.223 $-0.345$ +0.164 $-0.064$
Na@Rb +0.237 $-0.711$ +0.205 $-0.081$
Na@K@Rb -- -- +0.180 $-0.006$
Table 1: DFT data for atom pair exchanges. In the initial configurations, these clusters always contain a core with MC shells $(0,0)$ and $(0,1)$ made of Na atoms, while the other shells are made of K, Rb or K and Rb atoms. For MC@BG, the $(1,1)$ BG shell is added to the core, whereas for MC@BG@Ch1 a further $(1,2)$ shell is added (see Supplementary Figure 6a). For MC@MC, the $(0,2)$ shell is added to the core, whereas for MC@MC@MC a further $(0,3)$ shell is added. Exchanges of atom pairs are made on the initial configuration: a Na atom in the $(0,1)$ MC shell is exchanged with a neighbouring K/Rb atom of the outer shell (see Supplementary Figure 6b-e). The energy differences (in eV) between the configurations after and before the exchange of the atomic pair are reported, for the most favourable between the two possible inequivalent exchanges. The exchange is favourable for negative differences, unfavourable otherwise. Complete data are reported in Supplementary Tables 1-3.